# CURVALID: Geometrically-guided Adversarial Prompt Detection

### v1

@@ 4.2 TextCurv: Curvature at the word-level
To characterize the word-level geometric properties of benign and adversarial prompts, we analyze the curvature of word connections. The aim is to have curvature complement PromptLID by analyzing the word-level geometric properties of prompts, effectively identifying nearly all adversarial prompts. Specifically, curvature captures the relationships between words, revealing subtle semantic shifts based on word order, and uncovering local geometric differences between benign and adversarial prompts.
Previous research indicates that CNN activation creates a curved manifold, evidenced by the significantly higher intrinsic dimension estimated by Principal Component Analysis (PCA) compared to the GID estimated by TwoNN on activation data (Abdi & Williams, 2010; Facco et al., 2017; Ansuini et al., 2019). However, curvature differences between benign and adversarial prompts remain unexplored. Thus, we examine these differences in convolution layers as potential classification features.
Our goal is to establish a definition of text curvature based on existing mathematical definitions, with the curvature capturing semantic shifts according to word sequence and the strength of these shifts. Word sequence plays a crucial role in semantic analysis, helping to accurately capture the local geometric properties of prompts. We focus on the representations of prompts in the convolutional layers of the model $g$ as mentioned in Section 4.1, where the prompt data remains unflattened and in stacked lists of vectors at this stage, which can be viewed as word-level representation. Specifically, we extract the representations $\bm{z}_{2}$ and $\bm{z}_{3}$ from these convolutional layers for further analysis. This stage is critical, as it is where feature spaces are curved, according to prior research (Ansuini et al., 2019).
To capture semantic shifts between consecutive words in a prompt, we draw on Whewell’s equation, where the rate of directional change of a curve is represented by the tangential angular change. In NLP, this angular change is connected to the dot-product formula and cosine similarity, which indicate the semantic similarity or difference between two words (Mikolov, 2013; Levy et al., 2015). We assume that this theory also applies to modern word embeddings like GPT-2 and RoBERTa, and therefore, we define the rate of angular change in text curvature accordingly:
@@
[Text Curvature: Rate of angular change]
For any two consecutive word embeddings, denoted by $\vec{u}$ and $\vec{v}$, the rate of angular change, $d\theta$, is defined as:
@@
We focus on word vector magnitudes to capture the degree of semantic shift. Previous research suggests that magnitude reflects the semantic weight carried by each word and the tokenizers’ understanding of that word within its context (Schakel & Wilson, 2015; Reif et al., 2019). For example, common words tend to have smaller magnitudes due to their frequent use and limited semantic significance (Schakel & Wilson, 2015). Instead of summing vector norms to measure distance changes in curvature—which may seem intuitive and consistent with geometric principles—we sum the inverses of the vector norms. This approach is driven by the hypothesis that larger vector norms signify greater semantic importance, meaning that curvature should be inversely proportional to vector norms, capturing the larger semantic shifts between words.
Lemma 4.3.
@@
The rate of angular change in TextCurv is supported by Theorem 4.4, which links the angle $\theta$ to the difference in tangential angles. Further analysis of how word embedding vector norms relate to arc length is detailed in Appendix A.2.
@@
We employ the Gram-Schmidt process to construct an orthonormal basis for the tangent space, enabling us to express $\vec{u}$ and $\vec{v}$ as linear combinations of the basis vectors. In this basis, the angle $\theta$ between the vectors is determined using their inner product. We then compute the tangential angles of $\vec{u}$ and $\vec{v}$ from their projections onto the orthonormal basis. By isolating these tangential angles and subtracting them, we demonstrate that their difference is equivalent to $\theta$. The detailed proof is given in Appendix A.2. ∎
@@ 4.3 CurvaLID: Adversarial Prompt Classification by Curvature and LID
CurvaLID is an adversarial prompt detection method that filters out adversarial prompts before they are input to LLMs, ensuring their security. Since CurvaLID operates independently of LLMs, it provides a unified defensive performance across all LLMs. This differentiates it from existing SOTA defenses like input perturbation, prompt engineering, and adversarial training, which show varying performance across different adversarial prompts and LLMs. Moreover, CurvaLID’s evaluation is straightforward and standardized, avoiding the need for subjective human assessments or reliance on LLM judgments, which can raise robustness concerns (Chen et al., 2024; Raina et al., 2024).
CurvaLID involves four steps (see Figure 1, pseudo code in Appendix A.1). In the first step, we use our trained model $g$ (defined in Section 4.1) to classify different types of benign prompts, deriving the normal feature manifold. This is essential for amplifying the geometric and dimensional distinctions between benign and adversarial prompts in subsequent steps. In the second step, we extract the representations of both benign and adversarial prompts from $\bm{z}_{1}$, $\bm{z}_{2}$, and $\bm{z}_{3}$. Then, we compute PromptLID and TextCurv.
Then in the final step, we train a Multilayer Perceptron (MLP) to classify benign and adversarial prompts using the two mean TextCurv values, and the PromptLID as inputs. The MLP performs binary classification and filters out adversarial prompts before they reach the LLMs.
@@ 5 Experiments
This section highlights CurvaLID’s performance in classifying benign and adversarial prompts. We demonstrate the limitations of word-level LID due to stop words and punctuation. Lastly, we analyze TextCurv, emphasizing the difference between benign and adversarial prompts. Unless specified, the general experimental setup is as follows: we use a total of 2,519 testing prompts, comprising 1,200 benign prompts and 1,319 adversarial prompts. For the benign prompts, we randomly sampled 300 prompts from each of the Orca (Lian et al., 2023), MMLU (Hendrycks et al., 2020), AlpacaEval (Li et al., 2023a), and TruthfulQA (TQA) (Lin et al., 2021) datasets. The model $g$ is trained to classify these 4 benign datasets. Similarly, for the adversarial prompts, approximately 300 were randomly sampled from each of SAP (Deng et al., 2023a), DAN (Shen et al., 2023), MathAttack (Zhou et al., 2024b), and GCG (Zou et al., 2023). Detailed hyperparameter settings and configurations for the CNN and MLP are in Appendix A, B.1.4 and B.2.
We have included ablation studies and comparisons to other defenses in Appendix B.2 and B.4 respectively. It is important to note that SOTA defenses, such as input perturbation and adversarial training, are highly dependent on the specific adversarial prompts used and the particular LLMs targeted. In contrast, CurvaLID operates independently of the LLMs, providing consistent defensive performance across various models. This fundamental difference makes direct comparisons challenging. Nonetheless, we have attempted to summarize the performance of these methods and provide a comparative analysis with our method in Appendix B.4.
5.1 Adversarial Prompt Classification by CurvaLID
This section explores CurvaLID’s performance in classifying adversarial prompts, assessing its ability to identify mixed benign and adversarial inputs in real-world conditions. We also compare CurvaLID’s detection capabilities with existing defenses to demonstrate its effectiveness.
5.1.1 Classification on adversarial prompts
In this subsection, we demonstrate CurvaLID’s effectiveness in accurately classifying benign and adversarial prompts, including both English and non-English adversarial prompts. We compared a total of 8 datasets for our experiment, consisting of four benign datasets, namely Orca, MMLU, AlpacaEval, and TQA, and four adversarial datasets, namely SAP, DAN, MathAttack, and GCG. These datasets are the popular and SOTA English prompt datasets for testing the robustness of LLMs. From each dataset, we randomly selected at least 300 prompts, resulting in over 2,500 prompts for our classification experiment. We evaluated the ability of CurvaLID to distinguish between benign and adversarial prompts. The experimental setup used here is consistent with that of the following experiments unless specified. Further details about the datasets are in Appendix B.1.5.
CurvaLID on English adversarial prompts. The experimental results in Table 1 demonstrate CurvaLID’s high performance. The model achieved an overall accuracy of 0.992, with perfect accuracy of 1.00 (i.e., 100%) in identifying adversarial prompts and 0.984 accuracy in identifying benign prompts. Since adversarial prompts are detected before reaching the LLM, a detection accuracy of 1.00 directly translates to an attack success rate of 0, effectively nullifying adversarial attempts.
In addition to evaluating CurvaLID on the four main adversarial datasets, which contain a substantial number of prompts, we extended our experiments to other datasets like RandomSearch, AmpleGCG, Persuasive Attack, AutoDAN, and DrAttack (Chao et al., 2023; Andriushchenko et al., 2024; Liao & Sun, 2024; Zeng et al., 2024; Liu et al., 2023; Li et al., 2024). These datasets often include highly similar prompts or variants of those in our main experiments. CurvaLID can effectively identify all adversarial attacks with near 0.99 accuracy. Detailed results are in Appendix B.2.1.
CurvaLID on non-English adversarial prompts. We analyzed CurvaLID’s effectiveness in detecting non-English adversarial prompts by testing it on nine languages from the MultiJail dataset (Deng et al., 2023b). We randomly sampled 300 prompts from each of the 9 languages in the MultiJail dataset and tested them individually against 400 benign prompts, adjusted to avoid class imbalance. We gathered 100 prompts from each of four benign datasets randomly. Notably, retraining with benign prompts in different languages was unnecessary to maintain high accuracy. As shown in Table 2, CurvaLID achieved an overall accuracy and F1 score of 0.994, highlighting its robust ability to detect adversarial prompts across languages.
5.1.2 Analysis on CurvaLID
We conducted an extensive analysis of CurvaLID’s performance, including experiments with reduced training prompt sizes and training without adversarial prompts.
Impact of number of prompts on CurvaLID. We studied the impact of reducing the number of prompts on CurvaLID’s performance. We decreased the number of prompts in each dataset to 150 (from 300). Despite this reduction, CurvaLID maintained an overall accuracy of 0.988 and an F1 score of 0.988. Further details of the experimental results can be found in Appendix B.2.2.
Using CurvaLID in one-class classification problems. We experimented at modifying step 4 of CurvaLID by replacing the supervised MLP with unsupervised outlier detection methods such as the local outlier factor (LOF) (Breunig et al., 2000) or isolation forest (Liu et al., 2008), which do not require training on adversarial prompts. Despite the absence of adversarial examples during training, the LOF and isolation forest methods achieved a comparable detection accuracy of around 0.9. Therefore, our framework is task-independent and can be applied to various problem settings. Further details are provided in Appendix B.2.3.
@@ 5.1.2 Analysis on CurvaLID
Table 1: Performance metrics for CurvaLID on benign and adversarial datasets.
@@
Class Dataset Accuracy Class Acc. Overall Acc. F1 Benign Orca MMLU AlpEval TQA 0.984 0.992 0.992 0.968 1.000 0.983 0.986 Adv. SAP DAN MathAtk GCG 1.000 1.000 1.000 1.000 1.000
5.2 Analysis of local and global ID
@@
Table 3: Average TextCurv of benign and adversarial datasets across word embedding and CNN layers. Percentages in parentheses represent the increase in TextCurv for adversarial prompts compared to benign prompts.
@@
This subsection addresses the limitations associated with using token-level LID, for classifying benign and adversarial prompts. Adversarial prompts manipulate rarely encountered areas of the feature space using complex words and irregular combinations (Wallace et al., 2019; Ilyas et al., 2019; Ren et al., 2019). Existing work has shown that adversarial perturbation leads the representation to a subspace with distinct local dimensional properties (high LID) (Ma et al., 2018). Hence, we hypothesize that each word in adversarial prompts could lead to a representation with high LID. By utilizing token-level representations and aggregating the LIDs for each prompt by taking the average, it is possible to detect adversarial prompts.
Effectiveness of token-level LID values in separating benign and adversarial prompts. Our analysis revealed significant overlap in token-level LID values between benign and adversarial prompts. We calculated average word-level LID using RoBERTa embeddings, treating each word as a data point with its neighborhood defined by the prompt. As shown in Figures LABEL:fig:LID_SD and Appendix B.3.3, the average token-level LID across most datasets is around 10, making it difficult to differentiate between benign and adversarial prompts. Additionally, the standard deviations are notably high, with AlpacaEval exceeding 100 and DAN surpassing 200. This variability results in substantial overlap in LID values, reducing the effectiveness of token-level LID for distinguishing between prompts.
Investigating similarity benign and adversarial token-level LIDs. Next, we investigated the underlying reasons for the similarity in average token-level LID between benign and adversarial prompts. We focused on the average of the first 5 nearest-neighbor distances (NN-distance), which, as shown in Figure LABEL:fig:KNNdist, appear similar for both benign and adversarial datasets. We present the top 10 most common nearest-neighbors for different prompts in different datasets in Appendix B.3.2. It shows that common nearest-neighbors in the representation space were mostly stop words and punctuation. This suggests that word-level LID overlooks the sequential order of text and often relies on conjunctions, articles, and punctuation. These findings suggest that word-level LID is ineffective for detecting adversarial prompts.
Effectiveness of token-level LID after removing stop words and punctuation. To assess the impact of stop words and punctuation on LID, we recalculated word-level LID after removing them. As shown in Appendix B.3.3, this adjustment reduces the standard deviation to a maximum of 2.26, yielding more consistent LID values. However, the difference between benign and adversarial prompts’ LID remains insufficient for discrimination.
@@
Analysis on PromptLID. We analyze PromptLID across benign and adversarial datasets. As shown in Figure LABEL:fig:LID_comp, adversarial prompts exhibit a much higher average PromptLID (19.472) compared to benign prompts (3.509), highlighting its effectiveness in distinguishing adversarial prompts. The distribution of PromptLID between benign and adversarial prompts can be found in Appendix B.2.14.
Effectiveness of global intrinsic dimension in adversarial prompt detection. Global intrinsic dimension (GID) is another plausible approach for the word-level representation. It can avoid aggregating the LID for each word by assessing the GID for each word within the prompt and output a single value. We use the MLE-based estimate from Tulchinskii et al. (Tulchinskii et al., 2024). However, as shown in Figure LABEL:fig:GIDvsLength, GID shows no clear distinction between benign and adversarial datasets. Instead, it strongly correlates with prompt length, with Pearson and Spearman correlation coefficients of 0.92 and 0.98, respectively. Even after removing stop words and punctuation, the results were similar, with a Pearson correlation coefficient of 0.9, highlighting the limitations of GID in detecting adversarial prompts.
5.3 Curvature of benign and adversarial prompts
@@
We analyzed the average TextCurv of adversarial prompts in CurvaLID Step 2. To ensure TextCurv’s independence from the embedding model, we conducted curvature analysis using various embeddings, including GPT-2, BERT, XLNet, and DistilBERT (Radford et al., 2019; Devlin, 2018; Yang, 2019; Sanh, 2019). As shown in Table 3, adversarial prompts consistently exhibited at least 30% higher curvature than benign prompts across all embeddings. The TextCurv distribution for benign and adversarial prompts is also shown in Appendix B.2.14. These findings support our hypothesis that words in adversarial prompts are more irregular and complex than in benign prompts.
@@
We also demonstrated that CNN activation significantly amplifies the TextCurv differences between benign and adversarial prompts. Specifically, the mean TextCurv of adversarial prompts is at least 30% higher than that of benign prompts in the representations of both CNN layers. In contrast, when considering only the word embedding, the TextCurv for benign prompts based solely on word embedding is 4.91, compared to 5.42 for adversarial prompts. The difference in TextCurv is much smaller—just 13%, which is less than half of the difference observed in the CNN layers.

### v2

@@ 4.2 TextCurv: Curvature at the word-level
To characterise word-level geometric properties of benign and adversarial prompts, we analyse the curvature of word connections. The aim is to have curvature complement PromptLID by analyzing the word-level geometric properties of prompts, effectively identifying nearly all adversarial prompts. Specifically, curvature captures the relationships between words, revealing subtle semantic shifts based on word order, uncovering local geometric differences between benign and adversarial prompts.
Prior work shows that CNN activation creates a curved manifold, evidenced by the significantly higher intrinsic dimension estimated by Principal Component analysis (PCA) compared to the GID estimated by TwoNN on activation data (Abdi & Williams, 2010; Facco et al., 2017; Ansuini et al., 2019). However, curvature differences between benign and adversarial prompts remain unexplored. Thus, we examine these differences in convolution layers as potential classification features.
Our goal is to establish a definition of text curvature based on existing mathematical definitions, with the curvature capturing semantic shifts according to word sequence and the strength of these shifts. Word order plays a crucial role in semantic analysis, helping to accurately capture the local geometric properties of prompts. We focus on the representations of prompts in the convolutional layers of the model $g$ as mentioned in Section 4.1, where the prompt data remains unflattened and in stacked lists of vectors at this stage, which can be viewed as word-level representation. Specifically, we extract the representations $\bm{z}_{2}$ and $\bm{z}_{3}$ from these convolutional layers for further analysis. This stage is critical, as it is where feature spaces are curved, according to prior research (Ansuini et al., 2019).
To capture semantic shifts between consecutive words in a prompt, we draw on Whewell’s equation, where the rate of directional change of a curve is represented by the tangential angular change. In NLP, this angular change is connected to the dot-product formula and cosine similarity, which indicate the semantic similarity or difference between two words (Mikolov, 2013; Levy et al., 2015). We assume that this theory also applies to modern word embeddings like GPT-2 and RoBERTa, and therefore, we define the rate of angular change in text curvature accordingly.
@@
[Text Curvature: Rate of angular change] For any two consecutive word embeddings, denoted by $\vec{u}$ and $\vec{v}$, the rate of angular change, $d\theta$, is defined as:
@@
We focus on word vector magnitudes to capture the degree of semantic shift. Previous research suggests that magnitude reflects the semantic weight carried by each word and the tokenizers’ understanding of that word within context (Schakel & Wilson, 2015; Reif et al., 2019). For example, common words tend to have smaller magnitudes due to their frequent use and limited semantic significance (Schakel & Wilson, 2015). Instead of summing vector norms to measure distance changes in curvature, which may seem intuitive and consistent with geometric principles, we sum the inverses of the vector norms. This approach is driven by the hypothesis that larger vector norms signify greater semantic importance, meaning that curvature should be inversely proportional to vector norms, capturing larger semantic shifts between words.
Definition 4.3.
@@
The rate of angular change in TextCurv is supported by Theorem 4.4, which links $\theta$ to difference in tangential angles. analysis of how word embedding norms relate to arc length is in Appendix A.2.
@@
We apply the Gram–Schmidt process to form an orthonormal basis for the tangent space, expressing $\vec{u}$ and $\vec{v}$ as linear combinations of its vectors. In this basis, the angle $\theta$ follows from their inner product. We compute the tangential angles of $\vec{u}$ and $\vec{v}$ from their projections, and by subtracting them show the difference equals $\theta$. The full proof is in Appendix A.2. ∎
@@ 4.3 CurvaLID: Adversarial Prompt Classification by Curvature and LID
CurvaLID is an adversarial prompt detection method that filters out adversarial prompts before they are input to LLMs, ensuring their safety. Since CurvaLID operates independently of LLMs, it provides a unified defensive performance across all LLMs. This differentiates it from existing SOTA defences like input perturbation, prompt engineering, and adversarial training, which show varying performance across different adversarial prompts and LLMs. Moreover, CurvaLID’s evaluation is straightforward and standardised, avoiding the need for subjective human assessments or reliance on LLM judgments, which can raise robustness concerns (Chen et al., 2024; Raina et al., 2024).
CurvaLID involves four steps (see Figure 1, pseudo code in Appendix A.1). In Step 1, we use our trained model $g$ (defined in Section 4.1) to classify different types of benign prompts, deriving the normal feature manifold. This is essential for amplifying the geometric and dimensional distinctions between benign and adversarial prompts. In Step 2, we extract the representations of benign and adversarial prompts from $\bm{z}_{1}$, $\bm{z}_{2}$, and $\bm{z}_{3}$. In Step 3, we compute the PromptLID and TextCurv of each prompt using the representations from Step 2, capturing both sentence-level dimensionality and word-level curvature. In Step 4, we train a Multilayer Perceptron (MLP) to classify benign and adversarial prompts based on the two mean TextCurv values and the PromptLID. The MLP performs binary classification and filters out adversarial prompts before they reach the LLM.
@@ 5 Experiments
In our evaluation, we assessed both the reduction in attack success rate and the prompt classification accuracy of CurvaLID across a range of LLMs (Vicuna-7B-v1.1 (Chiang et al., 2023), LLaMA2-7B-Chat (Touvron et al., 2023), GPT-3.5 (Brown, 2020), PaLM2 (Anil et al., 2023), and Gemma-2-9B (Team et al., 2024)), and compared it against SOTA defences (SmoothLLM (Robey et al., 2023), Self-Reminder (Xie et al., 2023), Intentionanalysis (Zhang et al., 2024a), In-Context Demonstration (ICD) (Wei et al., 2023), RTT3d (Yung et al., 2024), and constrained SFT (Qi et al., 2025)). Our test set has 3,540 prompts, comprising 1,200 benign and 2,340 adversarial prompts. For benign data, we randomly sampled 300 from each of the Orca (Lian et al., 2023), MMLU (Hendrycks et al., 2020), AlpacaEval (Li et al., 2023b), and TruthfulQA (TQA) datasets (Lin et al., 2021). The model $g$ is trained to classify these four benign datasets. For adversarial data, approximately 300 were randomly sampled from each of SAP (Deng et al., 2023a), DAN (also known as the "In The Wild" dataset) (Shen et al., 2023), MathAttack (Zhou et al., 2024b), and GCG (Zou et al., 2023), while around 200 prompts were randomly selected from PAIR (Chao et al., 2023), RandomSearch (Andriushchenko et al., 2024), AmpleGCG (Liao & Sun, 2024), Persuasive Attack (Zeng et al., 2024), AutoDAN (Liu et al., 2023), and DrAttack (Li et al., 2024). Dataset details are in Appendix B.1.5. Unless specified otherwise, we use an 80/20 train–test split. CNN/MLP hyperparameters are in Appendix B.1.2 and B.1.4. Results are averaged over 10 runs for reliability, and ablations are reported in Appendix B.2.
5.1 Main Results
Table 1: We compare CurvaLID with SOTA defences. The best results are boldfaced. Results for other LLMs and defences are provided in Appendix B.4.1.
LLM defence GCG PAIR DAN AmpleGCG SAP MathAttack RandomSearch Vicuna-7B No defence 86.0 98.0 44.5 98.0 69.0 24.0 94.0 SmoothLLM 5.5 52.0 13.0 4.2 44.6 22.0 48.5 Self-Reminder 9.5 48.0 35.5 11.5 25.2 22.0 6.0 Intentionanalysis 0.0 8.5 3.3 0.3 0.23 20.0 0.0 ICD 0.2 5.2 40.4 0.9 32.8 22.0 0.2 RTT3d 0.2 0.3 22.0 3.5 33.5 20.2 2.5 CurvaLID 0.0 0.0 0.0 1.1 0.0 0.0 0.0 LLaMA2-7B No defence 12.5 19.0 2.0 81.0 9.5 11.7 90.0 SmoothLLM 0.0 11.0 0.2 0.2 1.2 11.2 0.0 Self-Reminder 0.0 8.0 0.3 0.0 0.0 11.1 0.0 Intentionanalysis 0.0 5.8 0.7 0.0 0.0 11.2 0.0 ICD 0.0 2.7 0.8 0.0 0.0 10.8 0.0 RTT3d 0.2 0.2 1.8 0.4 5.5 9.8 0.8 CurvaLID 0.0 0.0 0.0 0.0 0.0 0.0 0.0 PaLM2 No defence 14.9 98.0 49.7 88.9 55.1 18.9 91.9 SmoothLLM 5.5 38.7 6.7 7.2 41.2 9.8 45.3 Self-Reminder 2.3 36.7 22.3 4.7 21.4 13.3 3.7 Intentionanalysis 0.0 2.3 1.3 0.9 0.0 9.7 0.0 ICD 0.1 4.9 34.2 0.2 33.9 9.3 0.0 RTT3d 0.1 0.1 25.5 3.3 25.0 10.2 2.8 CurvaLID 0.0 0.0 0.0 0.0 0.0 0.0 0.0
@@ 5.1 Main Results
Table 2: Performance metrics for CurvaLID on benign and adversarial datasets. Standard deviations are shown in parentheses.
@@
Class Accuracy by Dataset Class Acc. Overall Acc. F1 Benign Orca MMLU AlpEval TQA 0.984 0.992 0.992 0.968 1.000 0.983 0.986 (0.012) (0.000) (0.008) (0.010) (0.009) Adv. SAP DAN MathAtk GCG 1.000 1.000 1.000 1.000 1.000 (0.000) (0.000) (0.000) (0.000) (0.000)
@@
Table 3: Average TextCurv of benign and adversarial datasets across word embedding and CNN layers. Percentage in parentheses shows the increase in TextCurv for adversarial prompts.
@@
(a) (a) Confusion matrix from CurvaLID on English adversarial prompts (corresponds to Table 3). (b) Average token-level LID for benign and adversarial prompts. Blue bars show LID for original prompts; red bars show LID after removing stopwords and punctuation. (c) Average nearest neighbor distances, blue for benign and red for adversarial prompts. (d) Average PromptLID for benign and adversarial prompts. Error bars in (b)–(d) show standard deviation over 10 runs.
Comparison of CurvaLID against baseline defences. We compared CurvaLID with five existing defences: SmoothLLM, Self-Reminder, Intentionanalysis, In-Context Demonstration (ICD), and RTT3d. Evaluations are conducted across four LLMs: Vicuna-7B-v1.1, LLaMA2-7B-Chat, GPT-3.5, and PaLM2, and against seven adversarial attacks, including GCG, PAIR, DAN, AmpleGCG, SAP, MathAttack, and RandomSearch. In addition, we compared CurvaLID with the constrained SFT defence (Qi et al., 2025), using the fine-tuned Gemma-2-9B model released by Qi et al. (2025). In our approach, if a prompt is classified as a jailbreak prompt, it is rejected. The comparison against SOTA defences across multiple LLMs, measured by ASR (%), is presented in Table 1. Appendix B.4.1 provides detailed experimental results and covers filters like Circuit Breakers (Zou et al., 2024), Llama Guard (Inan et al., 2023), and perplexity filtering (Alon & Kamfonas, 2023). Experimental result shows that CurvaLID effectively identifies adversarial prompts and rejects them before querying the LLM. Notably, CurvaLID is model-agnostic and performs consistently across LLMs, achieving superior results over baseline defences in most scenarios.
CurvaLID on adversarial prompts. The experimental results in Table 3 demonstrate CurvaLID’s high performance. The model achieved an overall accuracy of 0.992, with perfect accuracy of 1.00 (i.e., 100%) in identifying adversarial prompts and 0.984 accuracy in identifying benign prompts. Since adversarial prompts are detected before reaching the LLM, 1.00 accuracy implies a 0% attack success rate, effectively nullifying adversarial attempts. The corresponding confusion matrix is shown in Figure LABEL:fig:confusion_matrix. Notably, CurvaLID remains robust even when the number of prompts per dataset is halved to 150, achieving an accuracy and F1 score of 0.988 (see Appendix B.2.12).
In addition to the four main adversarial datasets, which contain a substantial number of prompts, we extended our experiments to PAIR, RandomSearch, AmpleGCG, Persuasive Attack, AutoDAN, DrAttack, and persona modulation attack. CurvaLID identified all adversarial attacks with near 0.99 accuracy (see Appendix B.2.1 and B.2.4). It also achieved over 0.9 accuracy on demonstration-based attacks, including In-Context Demonstration (Wei et al., 2023) and cipher-based attacks (Yuan et al., 2023) (see Appendix B.2.2), and 0.994 accuracy on adversarial prompts written in nine non-English languages (see Appendix B.2.3). We also evaluated CurvaLID on benchmarks, namely HarmBench adversarial prompts (Mazeika et al., 2024), achieving near-zero ASR, and over-refusal benchmarks (Cui et al., 2024; Röttger et al., 2023), where CurvaLID reduced harmful prompt acceptance by up to 30% (see Appendix B.2.5- B.2.6).
5.2 analysis on CurvaLID
We analyse CurvaLID under three settings: (1) training only on benign prompts, (2) comparing token-level LID with PromptLID, and (3) testing TextCurv across word embeddings.
Using CurvaLID in one-class classification problems. We modified step 4 of CurvaLID by replacing the supervised MLP with unsupervised outlier detection methods such as the local outlier factor (LOF) (Breunig et al., 2000) or isolation forest (Liu et al., 2008), which do not require training on adversarial prompts. Despite the absence of adversarial examples during training, LOF and isolation forest methods achieved comparable detection accuracy of around 0.9. Therefore, our framework is task-independent and can be applied to various problem settings. Full results are in Appendix B.2.13.
@@
Limitations of token-level LID. We provide the motivation for PromptLID by examining why traditional token-level LID fails to effectively distinguish benign from adversarial prompts. Adversarial inputs often manipulate rarely encountered regions of the feature space using complex words and irregular combinations (Wallace et al., 2019; Ilyas et al., 2019; Ren et al., 2019). Prior work has shown that such perturbations lead representations into subspaces with distinct local dimensional properties, typically exhibiting high LID (Ma et al., 2018). Based on this, we hypothesize that each word in an adversarial prompt may induce a representation with elevated LID, and that by aggregating these token-level values (e.g., via averaging), it might be possible to detect adversarial prompts.
However, our analysis reveals that this token-level approach, computed with RoBERTa embeddings and treating each word as a data point within its prompt-based neighbourhood, is ineffective at separating benign and adversarial inputs. As shown in Figures LABEL:fig:LID_SD and Appendix B.3.3, average token-level LID across datasets centers around 10 with high variability, causing substantial overlap between benign and adversarial classes. To investigate further, we analysed the first five nearest-neighbor distances, which also showed minimal differences between prompt types (Figure LABEL:fig:KNNdist). Appendix B.3.2 indicates the most common neighbors are stop words and punctuation, suggesting token-level LID is dominated by non-informative tokens and insensitive to sequential structure. Removing stop words and punctuation lowered the standard deviation to 2.26 (Appendix B.3.3), but the distinction between benign and adversarial prompts remained weak. These results reaffirm the limitations of token-level LID for adversarial detection and further motivate using PromptLID.
@@
We analyse PromptLID across benign and adversarial datasets. As shown in Figure LABEL:fig:LID_comp, adversarial prompts exhibit a much higher average PromptLID compared to benign prompts, highlighting its effectiveness in distinguishing adversarial prompts. The distribution of PromptLID between benign and adversarial prompts can be found in Appendix B.2.26.
@@
Generalisation to different word embedding. We analysed the average TextCurv of adversarial prompts in CurvaLID Step 2. To ensure TextCurv’s independence from the embedding model, we conducted curvature analysis using GPT-2 (Radford et al., 2019), BERT (Devlin, 2018), XLNet (Yang, 2019), and DistilBERT (Sanh, 2019). As shown in Table 3, adversarial prompts consistently exhibited at least 30% higher curvature than benign prompts across all embeddings. The TextCurv distribution for benign and adversarial prompts are in Appendix B.2.26. These findings support our hypothesis that words in adversarial prompts exhibit greater irregularity and complexity compared to those in benign prompts. More importantly, the results generalize across different embedding models, suggesting that adversarial and benign prompts differ fundamentally in their geometric properties.
@@
We also demonstrated that CNN activation significantly amplifies TextCurv differences between benign and adversarial prompts. The mean TextCurv of adversarial prompts is at least 30% higher than that of benign prompts in both CNN layers. In contrast, when using only the word embedding, the mean TextCurv is 4.91 for benign prompts and 5.42 for adversarial prompts, a much smaller difference of 13%, less than half of that observed in the CNN layers (see Appendix B.2.25).